HETAV PATEL

Grade 11,

There are three particles ABC which are situated at vertex of equilateral triangle ABCof side d at time T = 20 each of the particle moves with constant speed v always has its velocity along a b be always has be always has MC always has its velocity along CA at what time they will meet with each other

There are three particles ABC which are situated at vertex of equilateral triangle 
ABCof side d at time T = 20 each of the particle moves with constant speed v always has its velocity along a b be always has be always has MC always has its velocity along CA at what time they will meet with each other

Grade:11

1 Answers

Arun
5 years ago
Dear Arvind
 
 

velocity of A is v along AB 

the velocity of B is along BC . its component along BA is vcos60 = v/2. thus the separation AB decreases at the rate

 

v+v/2=3v/2

sice this rate is constt. the time taken in reducing the separation AB from d to 0 is 

 

t=d/(3v/2) =2d/3v